harleigh:

make tables to find the solution to 2^x=4^x-2 take the interger values of x between -4 and 4

Vocaloid:

you want to set up a table like so: 

Vocaloid:

for each row, evaluate the x value at each function for example, in row 1 column 2, you would evaluate 2^x for x = -4, and for row 1 column 3, you'd evaluate 4^x - 2 then keep going for x = -3, -2, etc. all the way up to 4 your solution is the x-value where columns 2 and 3 have the same value